OpenStudy (anonymous):

What does x² - 2xy + y² mean?

OpenStudy (dmnts):

\[(x-y)^{2}=x ^{2}-2xy+y ^{2}\]

OpenStudy (anonymous):

so can you type it out as (x-y) (x+y)?

OpenStudy (anonymous):

its.. (x - y)^2 = x^2 + y^2 - 2xy this is a formula..

OpenStudy (anonymous):

so can you type it out as (x-y) (x+y)?

OpenStudy (dmnts):

No, because\[(x-y)(x+y)=x ^{2}-y ^{2}\]

OpenStudy (anonymous):

then does (x-y)^2 simplify to (x-y)(x-y)?

OpenStudy (dmnts):

Yes

OpenStudy (unklerhaukus):

\[(x-y)^2\\=(x-y)(x-y)\\=x(x-y)-y(x-y)\\=x^2-xy-yx+y^2\\=x^2-2xy+y^2\]